fix) = x+2 X-2 using noo Tim faf AR)
Answer #1

Soeh f(x) = x + 2 lim (flerth) - f (x) es plantes - Blog xth-2 x-2 h lim (atht2 -x+2 ho lim Rath+2) (4-2) (x+2) (4th-2) ho [[lim x = 2yk *fx-2 ht 24-H -- he +244-24-2h ty] (x-2) (ath-2) h ho lim ho of .4k -4 (1-2) (x+h-2h (x-2)(x+0-2) = -4 (x-2) ² f